1. Assume that in 2015, the following prevails in the republic of Nurd

Y= $200

C= $160

S= $40

I (Planned)= $30

G= $0

T= $0


d. Suppose Y = $200, C = $160, S = $40, and I = $40. Is Nurd’s economy in equilibrium?


e. Starting with the situation in part d, suppose the govern-ment starts spending $30 each year with no taxation and continues to spend $30 every period. If I remains constant, what will happen to the equilibrium level of Nurd’s domestic product (Y)? What will the new levels of C and S be?


f. Starting with the situation in part d, suppose the govern-ment starts taxing the population $30 each year without spending anything and continues to tax at that rate every period. If I remains constant, what will happen to the equilibrium level of Nurd’s domestic product (Y)? What will be the new levels of C and S? How does your answer to part f differ from your answer to part e? Why?


1
Expert's answer
2021-12-15T11:36:18-0500

Solution:

d.). Yes, Nurd's economy is in equilibrium.

Y = C + I

200 = 160 + 40

200 = 200


e.). The equilibrium level of Nurd's domestic product (Y) will increase by 30.

Y = C + I + G

Y = 160 + 40 + 30 = 230

C will remain the same

New S = Y - C = 230 - 160 = 70


f.). The equilibrium level of Nurd's domestic product (Y) will decrease by 30.

Y = (C-t) + I = (160-30) + 40 = 130 + 40 =170

The new level of C = 160-30 = 130

The new level of S = Y - C = 170 - 130 = 40